OpenStudy (anonymous):

If a cup of tea has an initial temperature of 203 degree F in a room whose temperature is 70 degreeF, then according to Newton's law of cooling the temperature of the tea after t minute have passed is

OpenStudy (anonymous):

T (t) =70+133e^(-t/5) what is the average temperature of the tea one hour after the initial time?? please help!

OpenStudy (anonymous):

replace t by 1. \[T(1)=70+133e^{\frac{-1}{5}}\]

OpenStudy (anonymous):

i get 178.89 to two decimal places

OpenStudy (amistre64):

should 133 be 203?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

i see it now lol 203 - 70 right?

OpenStudy (anonymous):

the room temp is 70, 203-70=133 and it is the difference that decays

OpenStudy (anonymous):

how do you get that answer of 178.89 from that equation? b/c i got something different :(

OpenStudy (amistre64):

yeah, so t = -60 minutes if that is accurate....

OpenStudy (amistre64):

70 + (133 * (e^((-60) / 5))) = 70.0008172 is what google figures out

OpenStudy (amistre64):

or is this a recursion equation.... and not a particular solution?

OpenStudy (amistre64):

nah; particular it is; no place for a T{n-1} lol